OpenStudy (anonymous):

Please help medal+fanning!!! question down below Find the exact value of the radical expression in simplest form.

OpenStudy (anonymous):

\[c \sqrt{c}+12c \sqrt{c}-2c \sqrt{c}\]

OpenStudy (anonymous):

treat it like \[cx+12cx-2cx\] and combine like terms

OpenStudy (anonymous):

How do you combine like terms?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

I am attaching a blurb on combining like terms that may be of help.

OpenStudy (anonymous):

Thank you

Directrix (directrix):

These terms you are adding are all like terms because they have the same "last name." I'm going to write them so that will be clearer to see.

OpenStudy (anonymous):

How do I add up this one does the c become 1. If so should i add it like this 12+2+1

Directrix (directrix):

(1) * c√ c + (12)* c√ c - 2* (c√ c) = The common last name is (c√ c)

Directrix (directrix):

So you essentially have 1 + 12 - 2 = what and then add the last name of (c√ c) to that answer. Does that help?

OpenStudy (anonymous):

Ok so the answer is 11(csqrtc)

OpenStudy (anonymous):

\[11c \sqrt{c}\]

Directrix (directrix):

11c√ c

OpenStudy (anonymous):

"last name" i like that i will use it (steal it)

Directrix (directrix):

@satellite73 Consider it a gift. :)

OpenStudy (anonymous):

Loll you two are funny. Thanks so much directrix could you help check one more to see if I got this

Directrix (directrix):

Okay.

OpenStudy (anonymous):

Oh and was that the simplest form?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

11c√ c is the simplest radical form. Are you "into" rational exponents? Usually, the instructions for a problem tell you what form in which to express the answer.

Directrix (directrix):

Find the exact value of the radical expression in simplest form. Yes, it is in simplest radical form and it is exact, too. On the button.

OpenStudy (anonymous):

Loll ok thank you but what happens when the terms aren't so alike?

Directrix (directrix):

Then, you only combine the like terms.

OpenStudy (anonymous):

\[\sqrt{p^6}+\sqrt{9p^6}+\sqrt{3p^6}-\sqrt{p^2}\]

OpenStudy (anonymous):

What do I do with the unlike terms

Directrix (directrix):

First, simplify the square root of p^6, 9 p^6, and p^2. Those are perfect squares under the radical.

OpenStudy (anonymous):

13p^6

OpenStudy (anonymous):

minus p^2 is 13^4

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

What are you saying is the answer? I don't get what those posts mean in terms of the problem.

OpenStudy (anonymous):

Oh i thought the answer was 13^4 after combing the first like terms giving me f13p^6

Directrix (directrix):

I will need to write the problem by hand and begin the solution that way. Are you up for that? I don't care for LaTex myself.

OpenStudy (anonymous):

Loll sure thing but whats latex?

Directrix (directrix):

LaTex is whatever program you used to write what you see on the attachment.

OpenStudy (anonymous):

Oh the equation thing at the bottom below the comment I used

Directrix (directrix):

You can pick up from here. See the attachment. Keep in mind: In square roots, it takes two of the same thing under the radical to get one of them out. That is not precise mathematics but the idea is what counts.

OpenStudy (anonymous):

I'm a little confused on how you put the p3p3 instead of p6 @Directrix

OpenStudy (anonymous):

13 times p cubed times the square root of p − 2 times p times the square root of p 13 times p cubed times the square root of p − 2p 7p3 7p3 − p

Directrix (directrix):

You may need to review the exponent properties. p^3 times p^3 = p^6.

OpenStudy (anonymous):

I think it's a then

Directrix (directrix):

Here it is:

OpenStudy (anonymous):

Thats the final answer? @Directrix

Directrix (directrix):

That's what I got. Do you have questions? Sorry for the delay but I had to complete a Geometry proof I had helped another asker begin. But, he became impatient and signed off before I could finish. I finished anyway.

Directrix (directrix):

@HelloGoodmorning

OpenStudy (anonymous):

Lol No worries but do you see the four answers I put up one of those are the answers. 13 times p cubed times the square root of p − 2 times p times the square root of p 13 times p cubed times the square root of p − 2p 7p3 7p3 − p

Directrix (directrix):

Will you look at your text and check to be sure you LaTexed the third term correctly? I think that 3 is supposed to be a 9. I circled it on the atttachment.

OpenStudy (anonymous):

Oops sorry the one where it say \[\sqrt{3p^6}\] is suppose to say

OpenStudy (anonymous):

\[3\sqrt{p^6}\]

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

That's why the answer we got did not gee-haw with any of the options. Just a second.

Directrix (directrix):

@GodBlessYou Corrected problem and solution awaiting for your inspection and approval. :)

OpenStudy (anonymous):

Lol I'm so sorry but there's no need to solve it I found out what the answer was Thanks so much for coming when I need you

Directrix (directrix):

Okay, we can close out this thread now.

OpenStudy (anonymous):

Okay :D